Accounts receivable analysis The following data are taken from the financial statements of Sigmon Inc. Terms of all sales areShow the Excel Formulas Please.
Answer #1

Correct Answer:

20Y3

20Y2

Account receivable turnover ratio

8.4

7.3

Number of days sale in account receivables

43.6 Days

49.9 days

The collection of accounts receivable has increased. This can be seen in both, the Increase in accounts receivable turnover and the decrease in the collection period.

Working:

A

B

C=A/B

D=365/C

Year

Net Sales

Average Receivables

Account receivable turnover ratio

Number of days sale in account receivables

20Y3

$        56,96,000

$        6,80,000

8.4

43.6

20Y2

$        45,20,500

$        6,17,500

7.3

49.9

A

B

C=(A+B)/2

Year

Beginning account receivables

Ending Account Receivables

Average Receivables

20Y3

$           6,55,000

$        7,05,000

$        6,80,000

20Y2

$           5,80,000

$        6,55,000

$        6,17,500

End of answer.

Please give a thumbs-up, it will be highly appreciated.

Thanks.